OpenStudy (anonymous):

factor completely: see below:

OpenStudy (anonymous):

\[2a ^{2}-2ab+3ac-3bc\]

OpenStudy (mathmate):

Note the pattern 2,-2, 3,-3 which helps you to factor by grouping.

OpenStudy (anonymous):

its not coming to me

OpenStudy (anonymous):

its wrong

OpenStudy (anonymous):

i see 2 a 2 b 2 c's I dont get it

OpenStudy (mathmate):

What if we try: \(2a^2-2ab+3ac-3bc\\ =2a(a-b) + 3c(a-b) \\ \)

OpenStudy (anonymous):

Just regroup it again

OpenStudy (mathmate):

Factor out the common factor (a-b) and you'll have it.

OpenStudy (anonymous):

we have the same answer @mathmate

OpenStudy (anonymous):

ok common factor?? walk me through it please. how did you find common factor??

OpenStudy (mathmate):

I guess we should, especially in math! :)

OpenStudy (mathmate):

Something like this: 16-12 =4*4 - 4*3 =4(4-3)